# Multimodal Medical Code Tokenizer

### v2

@@ 3.2 Token Packing
Unlike image-text paired data, where modalities have significant overlap, the two modalities (text and graph) of medical codes used in this work are more distinct yet complementary: the text focuses on clinical definitions, while the graph encodes domain-specific relationships not fully conveyed through text alone. Therefore, in addition to capturing shared information, it is crucial to extract modality-specific information during the tokenization process to ensure that the resulting tokens are highly informative.
Inspired by (Wang et al., 2024a), we optimize the obtained tokens ($\mathcal{I}(\mathbf{e}_{t}^{s})$, $\mathcal{I}(\mathbf{e}_{g}^{s})$, $\mathcal{I}(\mathbf{e}_{t}^{c})$, $\mathcal{I}(\mathbf{e}_{g}^{c})$) by separately optimizing the shared and modality-specific information across these tokens and their corresponding quantized vectors.
For shared information, given the tokens $\mathcal{I}(\mathbf{e}_{t}^{c})$ and $\mathcal{I}(\mathbf{e}_{g}^{c})$, the objective first is to adopt Kullback–Leibler (KL) divergence optimize them by ensuring their distance matrices $\mathrm{dist}(\mathbf{e}_{t}^{c},\mathbf{C})$ and $\mathrm{dist}(\mathbf{e}_{g}^{c},\mathbf{C})$ follow a similar distribution, as following:
$\mathcal{L}_{\text{KL}}=D_{\text{KL}}(\text{softmax}(\mbox{-}\mathrm{dist}(\mathbf{e}_{t}^{c},\mathbf{C}))\parallel\text{softmax}(\mbox{-}\mathrm{dist}(\mathbf{e}_{g}^{c},\mathbf{C})))$
Then optimize the quantized vectors $\mathbf{\hat{e}}_{t}^{c}$, $\mathbf{\hat{e}}_{g}^{c}$ to be highly informative about the other modality, while minimizing redundancy, as follows:
@@
For specific information, given the tokens $\mathcal{I}(\mathbf{e}_{t}^{s})$ and $\mathcal{I}(\mathbf{e}_{g}^{s})$, the objective is to optimize these tokens by ensuring that the quantized vectors $\mathbf{\hat{e}}_{t}^{s}$, $\mathbf{\hat{e}}_{g}^{s}$ retain as much modality-specific information as possible, with minimal shared information, as follows:
@@
Based on the derivation of (Wang et al., 2024a), the common tokens could be optimized by the combination of InfoNCE loss between $\mathbf{\hat{e}}_{t}^{c}$ and $\mathbf{\hat{e}}_{g}^{c}$, and the alignment loss between $\mathbf{e}_{t}^{c}$ and $\mathbf{e}_{g}^{c}$. Then the loss for packing shared information across two modalities is formulated by: $\mathcal{L}_{token}^{c}=\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{t}^{c},\mathbf{\hat{e}}_{g}^{c})+\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{g}^{c},\mathbf{\hat{e}}_{t}^{c})-2\beta\mathbb{E}_{\mathbf{e}_{t}^{c},\mathbf{e}_{g}^{c}}(\mathbf{e}_{t}^{c}\cdot\mathbf{e}_{g}^{c})$. Additionally, the modality-specific tokens could be optimized by the combination of InfoNCE loss between $\mathbf{\hat{e}}_{t}^{c}$, $\mathbf{\hat{e}}_{g}^{c}$ and the quantized vectors of their augmented data, and the orthogonal loss between $\mathbf{\hat{e}}_{t}^{c}$, $\mathbf{\hat{e}}_{g}^{c}$, and $\mathbf{e}_{t}^{c}$, $\mathbf{e}_{g}^{c}$, respectively. Then the loss for packing specific information across two modalities is formulated by: $\mathcal{L}_{token}^{s}=\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{t}^{c},\mathbf{\tilde{e}}_{t}^{c})+\lambda\mathcal{L}_{\mathrm{orthogonal}}(\mathbf{\hat{e}}_{t}^{c},\mathbf{e}_{t}^{c})+\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{g}^{c},\mathbf{\tilde{e}}_{g}^{c})+\lambda\mathcal{L}_{\mathrm{orthogonal}}(\mathbf{\hat{e}}_{g}^{c},\mathbf{e}_{g}^{c})$.
Finally, we combine the shared and specific losses to form the overall token packing loss, $\mathcal{L}_{token}=\mathcal{L}_{token}^{c}+\mathcal{L}_{token}^{s}$, to optimize our obtained tokens, where $\beta$ and $\lambda$ are hyperparameters set to be equal. This approach enables MedTok to leverage both modality-shared and modality-specific information.
@@ 3.3 Training and Inference for MedTok
Training Stage. During the training stage, MedTok is trained by the sum of codebook loss $\mathcal{L}_{C}$, KL divergency loss $\mathcal{L}_{KL}$, loss for packing common and specific tokens across two modalities $\mathcal{L}_{token}$, where $\mathcal{L}=\mathcal{L}_{C}+\mathcal{L}_{KL}+\mathcal{L}_{token}$.
Inference Stage. After pre-training, MedTok can be seamlessly integrated into any model or pipeline dealing with medical codes, providing unified medical tokens for downstream tasks.
@@ 4 Experiments
Medical coding systems. We collected a total of 617,490 medical codes from eight commonly used coding systems: ICD-9 (Organization et al., 1988), ICD-10-CM (Fung et al., 2020), ICD-10-PCS (Averill et al., 2001), SNOMED CT (Donnelly et al., 2006), ATC (Miller & Britt, 1995), NDC (Palmer, 2006), CPT (Dotson, 2013), and RxNORM (Nelson et al., 2011), as shown in Table 1. These codes cover various events, including procedures, diagnoses, and medications. Each code is paired with a textual description from official documents and a subgraph from PrimeKG (Chandak et al., 2023). Details for data pre-processing are available in Appendix A.
@@
Baselines. To evaluate MedTok for tasks related to medical codes, we consider two types of tokenizers and five models based on EHR as baselines. The first type of tokenizer is text-based (e.g., bert-base-uncased (Devlin et al., 2018)), while the second is graph-based (e.g., VQGraph (Yang et al., 2024b)). Five EHR-based models are ETHOS (Renc et al., 2024), GT-BEHRT (Poulain & Beheshti, 2024), MulT-EHR (Chan et al., 2024), TransformEHR (Yang et al., 2023b), and BEHRT (Li et al., 2020). Details on implementation can be found in the Appendix B.
Evaluation setup. We consider two evaluation setups:
  • Out-patient evaluation: We evaluate MedTok together with patient prediction models on a dataset of patients who are not admitted to a hospital and consider two categories of tasks: 1 Operational Outcomes (OO), and 2 new diagnosis assignments (ND), following (Wornow et al., 2023). The OO includes MT, RA, and prolonged LOS. The new diagnosis assignments are used to predict the first diagnosis of a disease. Details of task definitions are in Appendix C.
4.1 MedTok tokenizer with in-patient EHR models
@@ 4.2 MedTok tokenizer with out-patient EHR models
Table 4 presents the AUPRC values for each baseline and its integration with MedTok across two task types on the out-patient HRShot dataset. The results reveal that our tokenizer has the most significant impact on mortality prediction in Operational Outcomes, achieving an average improvement of 11.30%. It also significantly improves the detection of new diagnoses of Acute MI, with an average improvement of 8.80%. As shown in Fig. 3, a comparison of three types of tokenizers further demonstrates the effectiveness of MedTok in integrating both graph and textual modalities. Additionally, when comparing performance across two in-patient datasets, we observe that MedTok is particularly beneficial for longitudinal data.
@@ 4.3 Ablation studies
In the ablation studies, to eliminate potential bias from different model architectures, we integrate MedTok with a vanilla Transformer-based model (e.g., TransformEHR) to examine the impact of the adopted modalities and the vocabulary size in MedTok on performance.
@@
Effects of modalities on MedTok. To evaluate the impact of the two modalities (text, graph) used in MedTok —medical code definitions and biological subgraphs derived from a biomedical knowledge graph—we assess its performance by removing the text and graph modalities separately. As shown in Fig. 4, MedTok, when leveraging both modalities, achieves the best performance across all tasks on three datasets. By comparing the performance of MedTok without the graph modality and MedTok without the text modality, we observe that both modalities contribute significantly to EHR-based prediction tasks. The graph modality benefits drug recommendation and new disease detection tasks, while the text modality proves essential for readmission prediction on MIMIC-III and operational outcomes in EHRShot. These findings emphasize the importance of incorporating the underlying information linked to medical codes.
@@
Effect of codebook size $N$. We further evaluate the impact of the codebook size on the performance of MedTok by training it with varying sizes and assessing its effectiveness across three distinct datasets integrated with TransformEHR. Fig. 5 presents the results for various codebook sizes across all tasks on the three datasets. The performance trends observed on MIMIC-III and MIMIC-IV are quite consistent, demonstrating a clear pattern where increasing the codebook size enhances the model’s performance. Specifically, the highest average performance is achieved when the codebook size is set to $N=12,000$, indicating that this size strikes an optimal balance between sufficient coverage of the medical vocabulary and avoiding overfitting.
@@
In contrast, when analyzing the performance on EHRShot, a dataset consisting of patients with longer visit histories than those in MIMIC-III and MIMIC-IV, we observe that MedTok benefits from a larger codebook size. For EHRShot, the highest average performance is achieved when the codebook size is increased to $N=21,000$. This suggests that for datasets with more extensive patient visit histories, a larger codebook may be more effective in capturing the underlying complexity of the medical information, thus improving the model’s predictive capabilities.
@@ 4.4 Using MedTok tokenizer for medical QA
MedTok demonstrates strong performance in EHR-based tasks, as shown in Tables 3-4. To further assess its capabilities, we explore its effectiveness in a generation task, specifically multiple-choice medical question answering (MedicalQA), where the goal is to select the correct answer to a given clinical question (Singhal et al., 2023). We evaluate whether MedTok enhances few-shot learning in MedicalQA by integrating its tokenized representations with a LLM (LLaMA3.1-8B (Dubey et al., 2024)). MedTok-generated tokens are used as prefix tokens, which provide structured medical context before the main input, allowing the LLM to incorporate additional domain knowledge.
For this evaluation, we use the MedDDx dataset (Su et al., 2024), which contains questions at three difficulty levels: Basic, Intermediate, and Expert. The process consists of three steps: (1) Disease code mapping – Extract disease mentions from each question and retrieve their corresponding medical codes. (2) Tokenization via MedTok – Convert medical codes into structured tokens using MedTok. (3) Prefix Token Fine-Tuning – Fine-tune LLaMA3.1-8B using MedTok tokens as prefix inputs before the question text. We fine-tune the model on 617 intermediate-level questions and evaluate performance on 227 expert-level and 158 basic-level questions. The results in Figure 5 show an accuracy improvement of 0.3% on MedDDx-Basic and 2.3% on MedDDx-Expert, suggesting that MedTok can enhance medical QA when used as a structured prefix representation.
Figure 5: A, The AUPRC values of MedTok with different codebook size $N$; B, The accuracy of LLaMA3.1-8B vs. MedTok +LLaMA3.1-8B on two MedDDx medical QA datasets.

### v3

@@ 3.2 Token packing
Unlike image-text paired data, where modalities have substantial overlap, the two modalities considered here (text and graph representations of medical codes) are more distinct but also highly complementary. Text provides the clinical definitions and describes each medical code in natural language. In contrast, the graph representation encodes domain-specific relationships such as disease co-occurrence, hierarchical groupings, and other medical ontologies. These relationships are not fully captured by text alone and introduce structured, expert-driven knowledge that is critical for many clinical and scientific applications. By considering both modalities, MedTok generates a representation that captures the cross-modality information and also preserves information that is unique to each modality. MedTok achieves this by extracting modality-specific features during the tokenization process, rather than relying on standard approaches that may blend or discard valuable distinctions. This ensures that MedTok’s tokens reflect both clinical language and the structured relationships present in the graph.
We pack tokens ($\mathcal{I}(\mathbf{e}_{t}^{s})$, $\mathcal{I}(\mathbf{e}_{g}^{s})$, $\mathcal{I}(\mathbf{e}_{t}^{c})$, $\mathcal{I}(\mathbf{e}_{g}^{c})$) by optimizing both modality-specific and cross-modality information between these tokens and their corresponding quantized vectors. To capture cross-modality information, we focus on the tokens $\mathcal{I}(\mathbf{e}_{t}^{c})$ and $\mathcal{I}(\mathbf{e}_{g}^{c})$. The objective uses Kullback-Leibler divergence to align their distance matrices, $\mathrm{dist}(\mathbf{e}_{t}^{c},\mathbf{C})$ and $\mathrm{dist}(\mathbf{e}_{g}^{c},\mathbf{C})$, such that they follow a similar distribution: $\mathcal{L}_{\text{KL}}=D_{\text{KL}}(\text{softmax}(\mbox{-}\mathrm{dist}(\mathbf{e}_{t}^{c},\mathbf{C}))\parallel\text{softmax}(\mbox{-}\mathrm{dist}(\mathbf{e}_{g}^{c},\mathbf{C})))$. Next, we optimize the quantized vectors $\mathbf{\hat{e}}_{t}^{c}$ and $\mathbf{\hat{e}}_{g}^{c}$ to maximize the information they carry about the other modality, while minimizing redundancy with their own modality. Specifically, we solve:
@@
For modality-specific information, MedTok optimizes tokens $\mathcal{I}(\mathbf{e}_{t}^{s})$ and $\mathcal{I}(\mathbf{e}_{g}^{s})$ by ensuring that the quantized vectors $\mathbf{\hat{e}}_{t}^{s}$ and $\mathbf{\hat{e}}_{g}^{s}$ retain maximal information about their respective modalities while minimizing cross-modality information between modalities. The optimal solutions are given by:
@@
Following Wang et al., we formulate the packing loss for cross-modality information as: $\mathcal{L}_{token}^{c}=\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{t}^{c},\mathbf{\hat{e}}_{g}^{c})+\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{g}^{c},\mathbf{\hat{e}}_{t}^{c})-2\beta\mathbb{E}_{\mathbf{e}_{t}^{c},\mathbf{e}_{g}^{c}}(\mathbf{e}_{t}^{c}\cdot\mathbf{e}_{g}^{c})$. Additionally, we formulate the packing loss for modality-specific information as: $\mathcal{L}_{token}^{s}=\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{t}^{c},\mathbf{\tilde{e}}_{t}^{c})+\lambda\mathcal{L}_{\mathrm{orthogonal}}(\mathbf{\hat{e}}_{t}^{c},\mathbf{e}_{t}^{c})+\mathcal{L}_{\mathrm{InfoNCE}}(\mathbf{\hat{e}}_{g}^{c},\mathbf{\tilde{e}}_{g}^{c})+\lambda\mathcal{L}_{\mathrm{orthogonal}}(\mathbf{\hat{e}}_{g}^{c},\mathbf{e}_{g}^{c})$.
We combine cross-modality and modality-specific losses into an overall token packing loss as: $\mathcal{L}_{token}=\mathcal{L}_{token}^{c}+\mathcal{L}_{token}^{s}$. We set hyperparameters $\beta$ and $\lambda$ to equal values throughout experiments in this study. This approach allows MedTok to leverage both cross-modality and modality-specific information.
@@ 3.3 Training and inference for MedTok
During the training stage, MedTok is trained by the sum of $codebook$ loss $\mathcal{L}_{vq}$, KL divergence loss $\mathcal{L}_{KL}$, token packing loss $\mathcal{L}_{token}$, where $\mathcal{L}=\mathcal{L}_{vq}+\mathcal{L}_{KL}+\mathcal{L}_{token}$. After pre-training, MedTok can be integrated into any model or pipeline dealing with medical codes, providing unified medical tokens for downstream tasks.
@@ 4 Experiments
Medical coding systems. We collected a total of 617,490 medical codes from eight commonly used coding systems: ICD-9 (Organization et al., 1988), ICD-10-CM (Fung et al., 2020), ICD-10-PCS (Averill et al., 2001), SNOMED CT (Donnelly et al., 2006), ATC (Miller & Britt, 1995), NDC (Palmer, 2006), CPT (Dotson, 2013), and RxNORM (Nelson et al., 2011), as shown in Table 1. These codes cover various events, including procedures, diagnoses, and medications. Each code is paired with a textual description from official documents and a subgraph from PrimeKG (Chandak et al., 2023). Details are available in Appendix A.
@@
Baselines. We consider two tokenizers and five EHR-based models based on EHR. The first type of tokenizer is text-based (e.g., bert-base-uncased (Devlin et al., 2018)), while the second is graph-based (e.g., VQGraph (Yang et al., 2024b)). Five EHR-based models are ETHOS (Renc et al., 2024), GT-BEHRT (Poulain & Beheshti, 2024), MulT-EHR (Chan et al., 2024), TransformEHR (Yang et al., 2023b), and BEHRT (Li et al., 2020). Implementation details are in the Appendix B.
Evaluation setup. We consider two evaluation setups:
  • Out-patient evaluation: We evaluate MedTok together with patient prediction models on a dataset of patients who are not admitted to a hospital and consider two categories of tasks: 1 Operational Outcomes (OO), and 2 new diagnosis assignments (ND), following (Wornow et al., 2023). The OO includes MT, RA, and prolonged LOS. The new diagnosis assignments are used to predict the first diagnosis of a disease. Details are in Appendix C.
4.1 MedTok tokenizer with in-patient EHR models
@@ 4.2 MedTok tokenizer with out-patient EHR models
Table 4 presents the AUPRC values for each baseline and its integration with MedTok across two task types on the out-patient EHRShot dataset. The results reveal that our tokenizer has the most significant impact on mortality prediction in Operational Outcomes, achieving an average improvement of 11.32%. It also significantly improves the detection of new diagnoses of Hyperlipidemia, with an average improvement of 6.00%. As shown in Figure 3, a comparison of three types of tokenizers further demonstrates the effectiveness of MedTok in integrating both graph and textual modalities. Additionally, when comparing performance across two in-patient datasets, we observe that MedTok is particularly beneficial for longitudinal data.
@@ 4.3 Ablation studies
To comprehensively understand the contributions of various components in MedTok, we conduct ablation studies on: (1) the effect of input modalities, and (2) the effect of shared and modality-specific optimization strategies. To ensure that our analysis is not confounded by architectural differences, we integrate MedTok with a standardized Transformer-based backbone (e.g., TransformEHR). This setup allows us to attribute performance differences directly to modalities and optimization strategies, rather than model architecture.
Multimodal learning in MedTok. To evaluate the impact of the two modalities (text, graph) used in MedTok —medical code definitions and biological subgraphs derived from a biomedical knowledge graph—we assess its performance by removing the text and graph modalities separately. As shown in Figure 4, MedTok, when leveraging both modalities, achieves the best performance across all tasks on three datasets. By comparing the performance of MedTok without the graph modality and MedTok without the text modality, we observe that both modalities contribute significantly to EHR-based prediction tasks. The graph modality benefits drug recommendation and new disease detection tasks, while the text modality proves essential for readmission prediction on MIMIC-III and operational outcomes in EHRShot. These findings emphasize the importance of incorporating the underlying information linked to medical codes.
@@
Effects of cross-modality and modality-specific information on MedTok. MedTok is built on two modalities, and we have analyzed the impact of each modality in Figure 4. The results show that the model performs best when using both modalities. Additionally, MedTok optimizes tokens by maximizing shared information between modalities while preserving modality-specific information. To assess the contribution of each loss component, we conducted ablation studies by retraining MedTok with different loss function combinations. We then apply the pretrained MedTok to all tasks across datasets to obtain its performances. The results (average AUPRC across all tasks) in Table 5 demonstrate that both shared and specific information optimization enhance performance, with the full optimization achieving the best results across all datasets. The vector quantization loss $\mathcal{L}_{vq}$ is the basic loss for tokenization. By optimizing both shared and specific information across two modalities, the performances are improved by 3.9%, 5.7%, and 9.1% on three datasets, respectively. The experimental results also show that shared and specific information contribute more to out-patient datasets.
Optimization MIMIC III MIMIC IV EHRShot
$\mathcal{L}_{vq}$ 0.373 0.387 0.287
$\mathcal{L}_{vq}$ + $\mathcal{L}_{token}^{c}$ + $\mathcal{L}_{KL}$ 0.379 0.409 0.314
$\mathcal{L}_{vq}$ + $\mathcal{L}_{token}^{s}$ 0.382 0.402 0.366
$\mathcal{L}$ 0.412 0.444 0.378
  • $\mathcal{L}=\mathcal{L}_{vq}+\mathcal{L}_{token}^{c}+\mathcal{L}_{KL}+\mathcal{L}_{token}^{s}$. $\mathcal{L}_{token}^{c}$ denotes the loss for packing cross-modality information, while $\mathcal{L}_{token}^{s}$ denotes the loss for packing modality-specific information.
Table 5: The averaged AUPRC across all tasks with different loss function combinations.
4.4 Hyperparameter sensitivity analysis
We next investigate the impact of hyperparameters on MedTok’s performance across all datasets. MedTok is trained with three key hyperparameters: two weighting coefficients, $\lambda$ and $\beta$, which control the contribution of shared and specific information loss components, and a $codebook$ size parameter, $N$, which determines the number of discrete tokens available for representation.
Effects of loss weight $\lambda$ and $\beta$. To ensure that MedTok treats shared and specific information equally, we set $\lambda=\beta$, where $\lambda$ and $\beta$ are the weighting coefficients for the shared and specific information loss terms, respectively. The average AUPRC scores across all tasks are presented in below Figure 5A. The results demonstrate the influence of hyperparameter choices on model performance across the three datasets. Based on our findings, we recommend setting $\lambda=\beta=0.1$ for in-patient settings and $\lambda=\beta=0.01$ for out-patient settings.
Figure 5: A, The AUPRC values of MedTok with different weighting coefficients $\lambda,\beta$; B, The AUPRC values of MedTok with different $codebook$ size $N$.
@@ 4.4 Hyperparameter sensitivity analysis
Effects of $codebook$ size $N$. We further evaluate the impact of the $codebook$ size on the performance of MedTok by training it with varying sizes and assessing its effectiveness across three distinct datasets integrated with TransformEHR. Figure 5B presents the results for various $codebook$ sizes across all tasks on the three datasets. The performance trends observed on MIMIC-III and MIMIC-IV are quite consistent, demonstrating a clear pattern where increasing the $codebook$ size enhances the model’s performance. Specifically, the most stable average performance is achieved when the $codebook$ size is set to $N=12,000$, indicating that this size strikes an optimal balance between sufficient coverage of the medical vocabulary and avoiding overfitting.
@@
In contrast, when analyzing the performance on EHRShot, a dataset consisting of patients with longer visit histories than those in MIMIC-III and MIMIC-IV, we observe that MedTok benefits from a larger $codebook$ size. For EHRShot, the highest average performance is achieved when the $codebook$ size is increased to $N=24,000$. This suggests that for datasets with more extensive patient visit histories, a larger $codebook$ may be more effective in capturing the underlying complexity of the medical information, thus improving the model’s predictive capabilities.
Figure 6: The accuracy of LLMs vs. MedTok+LLMs on three medical QA datasets.
@@ 4.5 Using MedTok tokenizer for medical QA
MedTok demonstrates strong performance in EHR-based tasks, as shown in Tables 3-4. To further assess its capabilities, we explore its effectiveness in a generation task, specifically multiple-choice medical question answering (MedicalQA), where the goal is to select the correct answer to a given clinical question (Singhal et al., 2023).
We evaluate whether MedTok enhances few-shot learning in MedicalQA by integrating its tokenized representations with three LLMs (LLaMA3.1-8B (Dubey et al., 2024), Qwen2.5-7B (Hui et al., 2024), MMedLM (Qiu et al., 2024)). MedTok-generated tokens are used as prefix tokens, which provide structured medical context before the main input, allowing the LLM to incorporate medical codes.
For this evaluation, we use three medical QA datasets, including MMLU (Hendrycks et al., 2021), MedDDx (Su et al., 2024), and AfrimedQA (Olatunji et al., 2024). In addition, MedDDx dataset contains questions at three difficulty levels: Basic, Intermediate, and Expert.
The process consists of three steps: (1) Disease code mapping – Extract disease mentions from each question and retrieve their corresponding medical codes. (2) Tokenization via MedTok – Convert medical codes into structured tokens using MedTok. (3) Prefix token fine-tuning – Fine-tune LLMs using MedTok tokens as prefix inputs before the question text. We fine-tune the model on the MedMCQA dataset (Pal et al., 2022), and then evaluate the fine-tuned model on MMLU, MedDDx, and AfrimedQA.
The results in Figure 6 show an accuracy improvement across all datasets compared with LLMs, suggesting that MedTok can enhance medical QA when used as a structured representation of medical codes and integrated with an LLM through prefix tuning.
4.6 Interpretability of MedTok
To this end, we selected a subset of patients classified as high risk for Hyperlipidemia by MedTok+TransformEHR, where these patients have no records of Hyperlipidemia before. We then counted the tokens assigned to these patients and identified those appearing more than 100 times, as shown in Figure 7. We then mapped these frequent tokens to medical codes, with the most overlapping codes being Rosuvastatin 5 mg Oral Tablet (RxNorm 2669980), Burn of skin (SNOMED CT 147087003), Type 2 diabetes mellitus without complication (disorder) (SNOMED CT 373555004), and Hyperlipidemia (SNOMED CT 285605009). They are closely related to Hyperlipidemia. Rosuvastatin corresponds to medications commonly prescribed for lipid disorders. The other three medical codes represent clinical diagnoses or findings associated with hyperlipidemia-related cardiovascular risk. It suggests that MedTok captures key medical concepts related to Hyperlipidemia, supporting its predictive capability.
Figure 7: Top 100 frequent token IDs appearing in patients at high risk of Hyperlipidemia.